A man is looking at the sky passing an airplane at a height of 7200 m and just when it is above his head he observes that a man falls as observed in the closed circle. Determine the time it takes for the falling man to impact the ground? (g = 9.8 m / s2), do not consider air friction.
Answer #1

The time taken to reach the ground can be calculated by using the equation,

s = ut + \frac{1}{2}at^2 \Rightarrow t = \sqrt{\frac{2s}{a}}

Substiuting values we can write,

\Rightarrow t = \sqrt{\frac{2(7200)}{9.8}} = 38.33 s